Diffusion Morning Report 04/26/12

Ted Berg tberg at stanford.edu
Thu Apr 26 06:47:20 PDT 2012


Good Morning All,
     Looking better this morning. New items: Tylan 1 reported low growth 
rate.
     Leftovers: POCl3 and BBr3 sensor issues, Epi Silane MFC, Tylan Burn 
box, Thermco 3 and 4 TLC issues, Tylan 2 poor uniformity
     Resolved: Dopant and lift issues on Epi resolved, Working on Burn 
box solution.

					
	Date: 	4/26/12 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*yellow* 	*low growth rate* 	** 	** 	*3*
Tylan 2 	*yellow* 	*bad uniformity* 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*yellow* 	*tgo sensor issue-ran BBr3 run-observed* 	** 
** 	**
Tylan 6-POCl3 	*yellow* 	*tgo sensor issue tube reinstalled-ran POCl3 
run* 	** 	** 	**
Tylan 7-TEOS2 	*down* 	*burn box* 	*14.39/30* 	** 	**
Tylan 8-SiGe 	*down* 	*re-assembling-pump out again* 	*3.4/30* 	** 	**
Tylan 9-Poly 	*down* 	*burn box* 	*15.78/30* 	** 	*1*
Tylan 10-Nitride 	*down* 	*burn box* 	*14.22/30* 	** 	*1*
Tylan 11-LTO 	*down* 	*burn box* 	*0./30* 	** 	*1*
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*up* 	** 	** 	** 	**
Thermco 3 	*yellow* 	*testing TLC* 	** 	** 	**
Thermco4 	*yellow* 	*bad source controller-no TLC* 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	*
* 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*yellow* 	*needs test* 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi Cham. A 	*yellow* 	*SiH4 issues* 	** 	** 	*2*
AMAT Epi Cham. B 	*down* 	** 	** 	** 	**
TEL Spa 	*yellow* 	*needs test* 	** 	** 	**
Burn Box 	*down* 	*high pressure* 	** 	** 	**
SEM 4160 	*up* 	** 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	*on* 	*crossover ok* 	*changeover working well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20120426/1dadf6d8/attachment.html>


More information about the diffusion mailing list